OpenStudy (anonymous):

Need Help An expression is shown below: 2x3y + 18xy - 10x2y - 90y Part A: Rewrite the expression so that the GCF is factored completely. (3 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (4 points) Part C: If the two middle terms were switched so that the expression became 2x3y - 10x2y + 18xy - 90y, would the factored expression no longer be equivalent to your answer in part B? Explain your reasoning. (3 points)

OpenStudy (anonymous):

I think this is the right answer for part 1 but im not sure 2x^3y + 18xy - 10x^2y - 90y= 2y(x^3-5x^2+9x-45)

OpenStudy (er.mohd.amir):

right

OpenStudy (anonymous):

i dont understand how to do part 2

OpenStudy (er.mohd.amir):

2xy(x2+9)-10y(x2+9) =(2xy-10y)(x2+9) =2y(x-5((x2+9) =

OpenStudy (er.mohd.amir):

2y (x-5) (x2+9)

OpenStudy (anonymous):

for part 3 would it be the same answer for part 2 but switched around?

OpenStudy (er.mohd.amir):

yes ans is same since by associative property.

OpenStudy (anonymous):

thx for the help